OpenStudy (anonymous):

@ amistre64 , using the method that you taught me, how would I solve: A cyclist travels 20km in 4hrs. What speed did the cyclist cycle at?

OpenStudy (anonymous):

You need to view this as a units problem. You want a solution in units of \[\frac{km}{hr}\] You have 20km and 4hrs. What do you think you need to do to get the units we seek?

OpenStudy (saifoo.khan):

\[speed = \frac{20}{4}\] \[speed = 5km/hr\]

OpenStudy (anonymous):

Keep in mind the value sticks with the unit: \[\frac{20km}{4hr}\]

OpenStudy (anonymous):

sorry everybody keeps confusing me with their different methods

OpenStudy (anonymous):

At the end of the day you stick with the method that works for you. The answer should be the same regardless of method.

OpenStudy (anonymous):

As Saifoo stated above you shoudl get a result of \[5\frac{km}{hr}\]

OpenStudy (anonymous):

http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e5e651e0b8b1f45b498e7d6, this is the method I would like to use if possible

OpenStudy (anonymous):

the method I was using is similar to cross multiplication

OpenStudy (anonymous):

I looked at http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e5e651e0b8b1f45b498e7d6 The reason why amistre64 used crossmultiplication was to convert mintues into hours. In this problem you already have the hours. No need to convert.

OpenStudy (anonymous):

bloody hell

OpenStudy (anonymous):

so I cant use it universally

OpenStudy (anonymous):

You can use what Amistre showed you. He was cancelling units. i.e. converting minutes to hours: \[30 minutes\frac{1 hours}{60 minutes}=\frac{1}{2} hours\] We just didn't need to cancel units.